RC Meeting, Wednesday, October 13 , 2021

From HallCWiki
Jump to: navigation, search


RC Meeting Information

RC Daily Meetings are at 10 am, remote https://bluejeans.com/861439920

Accelerator Status

  1. Beyond usual trips and machine hiccups we have received beam fairly smoothly, even with the hall A beamline being set up over night
  2. Setup for Hall A will continue on 1pass for at least 2 days
    • Their production data starts with 3pass so we will hear from them soon when they will want to do a pass change

Hall C Status

  1. The hall is hanging in there and experiment is in good shape
  2. So far we have collected approximately 760 pions in this configuration
  3. Good luck to Dave Gaskell in his double role as RC and spokesperson over the next week

Current Run Plan

  1. Continue data taking until early Nov

Current Issues

  1. Replace flow switches for magnets: next opportunistic access (Joe Beaufait) ( PENDING )
  2. The HB has seen some fluctuations in the LHe levels. At the next opportunistic access Heidi would like to investigate the cryo line. ( PENDING )
    • Heidi believes this is related to a vacuum issue and the access is for diagnostic purposes only. Can be escorted during regular office hours (do not call her in at midnight for this)
    • If indeed it proves to be the case the only remedy would be to warm up the magent and pump the region (which would be a multiple day operation)
    • So far the PID loops have done their job and Heidi (together with the shift crew) are keeping a close eye on the LHe levels
  3. Water flow to the dump aperture needs to be replaced: needs 4+h and Steve Suhring believes that it can wait until the SAD ( PENDING )
  4. Target ladder camera was identified to be bad: Joe Beaufait will purchase a replacement in the next funding cycle ( PENDING )
    • Still need to follow-up with Joe about when the purchase can be done.

Special Requests

Useful Links